# A detailed dive into fitting strategies for GRB afterglows with contamination: A case study with kilonovae

### v1

@@ 5 Conclusions
Detailed observations are beginning to readily reveal contamination of a GRB afterglow with additional components. In this work, we have explored several strategies for fitting such observations and inferring the properties of the individual afterglow and additional component. In particular, we have focused on the scenario of a kilonova and an afterglow motivated by recent observations of long gamma-ray bursts afterglows with infrared excesses consistent with a kilonova (Rastinejad et al., 2022; Levan et al., 2023; Yang et al., 2024).
@@
We first determine the region of parameter space (time, wavelength and afterglow and kilonova properties) where contamination is more likely. In general, kilonova emission is brighter and lasts longer in the lower frequency NIR bands, up to post GRB, in comparison with higher frequency UV bands where the onset of decay can begin as soon as post GRB. In an on-axis contamination scenario, the kilonova will be observed as a bump in the lightcurve, the significance of which is determined by the flux density of the afterglow component. The optimal observation window in this case is approximately . In an off-axis scenario, the kilonova will always initially dominate the signal, which is favourable for kilonova parameter estimation. The takeaway, regardless of viewing angle, is to observe as early as possible and make use of NIR bands if aiming to detect a kilonova, which is consistent with current observational strategy.
@@
We then explored different strategies to fit the data and estimate the parameters of the kilonova and afterglow using synthetic, simulated data where we knew the true input. In many cases where the afterglow and kilonova are fit individually, a predicted lightcurve may look reasonable to the eye but the estimated parameters are completely inconsistent with the true input, highlighting the dangers of this strategy for correct interpretations. By investigating different contamination scenarios, we find that for on-axis observations, the kilonova cannot be recovered correctly using individual model fitting. The more commonly used strategy of subtracting a best-fit afterglow and then fitting a kilonova fares better, but again the kilonova parameters can be biased. We note that this also does not produce a fully marginalised posterior, but instead a posterior on the kilonova parameters conditioned on the afterglow model and best-fit parameters, which is not ideal. We find that the best fitting strategy is to jointly fit both the afterglow and kilonova and have showcased a Bayesian framework (publicly available via Redback) to perform such analyses.
@@
When the afterglow is missing observations in the radio and X-ray bands, it can be difficult to distinguish between data points that are strictly from afterglow contributions and those that are strictly from kilonova contributions. This is especially true for off-axis systems where it is possible for data points from the initially kilonova dominated signal to be fit as if they were an extension of the afterglow. Individual parameter inference is unreliable in these scenarios and is difficult to improve via a subtraction technique due to contradicting fits between the afterglow and kilonova. The only solution here is to again use a joint-fitting approach. However, while X-ray and radio data are not necessary to ensure a successful joint fit, these observations hold the power to decrease computational time by improving sampling efficiency and significantly improve the precision of predictions and, therefore, should remain a priority when observing.
@@
In this paper, we have demonstrated that where there is possible contamination between a kilonova and an afterglow, the use of a joint model for simultaneous transient inference is the ideal methodology for accurate and robust parameter estimation. Joint fitting also proves to be the most reliable method for inference, succeeding in situations where individual fits or subtraction techniques fail. While the combined model has been successful on our synthetic data sets, we recognise that there is significant room for development to make the combined model more realistic. In particular, our combined model implicitly assumes that the afterglow and kilonova emission is uncoupled, contrary to numerical simulations that show that kilonova ejecta properties change due to the traversal of a jet (Klion et al., 2021; Nativi et al., 2021). Furthermore, our analytic models, are limited by the assumption of a constant, grey opacity, which implicitly builds a shape to the overall kilonova lightcurve that numerical simulations of kilonova lightcurve do not always obey. An alternative approach to the semi-analytical models we use would be to use machine-learning based surrogates to model the numerical simulations of kilonovae (e.g., Lukošiute et al., 2022). However, these simulations are performed under the assumptions of local thermal equilibrium (LTE), which is invalid for most kilonova parameters past (Pognan et al., 2023), and can have a significant impact on the lightcurves past this epoch. Few studies have performed non-LTE kilonova modelling due to dramatically higher computational cost and nuclear data required to compute such lightcurves. As this work uses only simulated data, non-LTE effects are not considered, and the simplified semi-analytical formulae for kilonovae are considered the ‘true’ input for our analysis, even past , which could affect our conclusions. Further detailed modelling may also account for an afterglow jet with angular structure. Both such changes may help with the accuracy of parameter estimation, but it is unlikely that they will significantly change our conclusions based on the choice of methodology.
@@
An important caveat to our exploration of the idealised approach to fitting is the consideration that the models themselves are also inherently uncertain. And so while the choice of fitting strategy, e.g. subtraction may bias our results, the bias is (at least in most of our simulated scenarios) smaller than the systematic uncertainty in the kilonova models themselves due to uncertain underlying physics (Sarin & Rosswog, 2024; Brethauer et al., 2024). Therefore, we expect broad (order of magnitude estimates) of kilonova properties from subtraction-based fitting strategies (e.g., Rastinejad et al., 2022; Rastinejad et al., 2024) are likely robust. However, the bias is likely more of a concern for the relatively more certain afterglow parameters. We leave a detailed exploration with real observations for future work. We note that regardless of the size of bias a fit based on individual component fitting or subtraction creates, it is not the ideal quantity one desires when estimating parameters through Bayesian inference; a marginalised posterior distribution, dependent only on the observed data. Instead these alternative approaches implicitly condition the posterior on the choice and parameters of a best-fit afterglow model or some other underlying assumption. This is not ideal when what we truly desire is a posterior marginalising over every other degree of freedom.
@@ Acknowledgements
W. Wallace would like to thank Nordita for hosting her visit for the duration of this project, and acknowledges support for the mobility from a Turing Scheme grant funded by the UK government and awarded through the University of Bath. Through the CRESST II CoOp W.Wallace is also supported by the National Aeronautics and Space Administration (NASA) under award number 80GSFC21M0002. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of NASA. N.S acknowledges support from the Knut and Alice Wallenberg foundation through the “Gravity Meets Light” project, and the support through the Nordita Fellowship. Nordita is supported in part by NordForsk.

### v2

@@ 5 Conclusions
Detailed observations are beginning to readily reveal contamination of a GRB afterglow with additional components. In this work, we have explored several strategies for fitting such observations and inferring the properties of the individual afterglow and additional component. In particular, we have focused on the scenario of a kilonova and an afterglow motivated by recent observations of long GRB afterglows with infrared excesses consistent with a kilonova (Rastinejad et al., 2022; Levan et al., 2023; Yang et al., 2024).
@@
We first determine the region of parameter space (time, wavelength and afterglow and kilonova properties) where contamination is more likely. In general, kilonova emission is brighter and lasts longer in the lower frequency NIR bands, up to post GRB, in comparison with higher frequency UV bands where the onset of decay can begin as soon as post GRB. In an on-axis contamination scenario, the kilonova will be observed as a bump in the lightcurve, the significance of which is sensitive to the spectral luminosity of the afterglow component. The optimal observation window in this case is approximately . In an off-axis scenario, the kilonova will always initially dominate the signal, which we have shown is generally favourable for kilonova parameter estimation. The takeaway, regardless of viewing angle, is to observe as early as possible and make use of NIR bands if aiming to detect a kilonova, which is consistent with current observational strategy.
@@
We then explored different strategies to fit the data and estimate the parameters of the kilonova and afterglow using synthetic, simulated data where we know the true input. We apply the different strategies to different contamination scenarios, and account for changes in posterior distributions due to different random noise and data realizations within the same scenario. In many cases where the afterglow and kilonova are fit individually, we find that a predicted light curve may look reasonable to the eye while the estimated parameters are inconsistent with the true input, highlighting the dangers of this strategy for correct interpretations. In our on-axis observation scenarios, the inferred kilonova cannot be recovered using individual model fitting, and we find this holds true across different parameter variations and data realisations. For off-axis observations, the kilonova is generally fit better due to its dominance in optical bandpasses at early times, but exploring different parameter combinations shows that this does not rule out the possibility of encountering severe biases with an individual model inference approach. The more commonly used strategy of subtracting a best-fitting afterglow and then fitting a kilonova fares better, but again the kilonova parameters can be biased and in some cases can even get worse. We note that this also does not produce a fully marginalised posterior, but instead a posterior on the kilonova parameters conditioned on the afterglow model and best-fitting parameters, which is not ideal. We find that the best strategy across all observation angles is to jointly fit both the afterglow and kilonova, and have showcased a Bayesian framework (publicly available via Redback) to perform such joint analyses. The resulting inferred values using the joint method can be at worst comparable to individual methods and at best better than 30 more accurate.
@@
When the afterglow is missing observations in the radio and X-ray bands, it can be difficult to distinguish between data points that are strictly from afterglow contributions and those that are strictly from kilonova contributions. This is especially true for off-axis systems where data points from the initial kilonova-dominated signal can be fit as if they were an extension of the afterglow. Individual parameter inference is unreliable in these scenarios and is difficult to improve via a subtraction technique due to contradicting fits between the afterglow and kilonova. The only solution here is to again use a joint-fitting approach. However, while X-ray and radio data are not necessary to ensure a successful joint fit, these observations hold the power to decrease computational time by improving sampling efficiency and significantly improve the precision of predictions and, therefore, should remain a priority when observing.
In the limiting case where an afterglow is observed but there is no true kilonova contamination, we find that the independent afterglow fit is the only method that will reliably recover the correct parameters due to biases induced from a model misspecification, highlighting the need to perform posterior predictive checks that can test for biases (e.g., Gelman et al., 1996). The afterglow properties from a joint fit are more likely to have an accurate (biased) recovery when the data is less (more) informative within the window where kilonova observation is favourable. Likewise, when the data is less (more) informative in this regime, the joint and individual kilonova fits can predict a transient that appears (un)reasonable. Kilonova fits that do not appear sensible, including those resulting from afterglow subtraction when very few data points remain, often have posteriors that strongly rail against the priors and may serve as an indicator of a misspecified model. We note that when data is of low quality, caution should be taken when interpreting the results of a joint fit that predicts a kilonova, even one that appears to have reasonable parameters.
@@
In this paper, we have demonstrated that where there is contamination between a kilonova and an afterglow, using a joint model for simultaneous transient inference is the ideal methodology for accurate and robust parameter estimation. Joint fitting also proves to be the most reliable method for inference, succeeding in situations where individual fits or subtraction techniques fail. While the combined model has been successful on our synthetic data sets, we recognise that there is significant room for development to make the combined model more realistic. In particular, our combined model implicitly assumes that the afterglow and kilonova emission is uncoupled, contrary to numerical simulations that show that kilonova ejecta properties change due to the traversal of a jet (Klion et al., 2021; Nativi et al., 2021). Furthermore, our analytic models are limited by the assumption of a constant, grey opacity, which implicitly builds a shape to the overall kilonova light curve that numerical simulations of kilonova light curve do not always obey. An alternative approach to the semi-analytical models we use would be to use machine-learning based surrogates to model the numerical simulations of kilonovae (e.g., Lukošiute et al., 2022). However, these simulations are performed under the assumptions of local thermal equilibrium (LTE), which is invalid for most kilonova parameters past (Pognan et al., 2023), and can have a significant impact on the light curves past this epoch. Few studies have performed non-LTE kilonova modelling due to dramatically higher computational cost and nuclear data required to compute such light curves. As this work uses only simulated data, non-LTE effects are not considered, and the simplified semi-analytical formulae for kilonovae are considered the ‘true’ input for our analysis, even past , which could affect our conclusions. Further detailed modelling may also account for an afterglow jet with angular structure. Both such changes may help with the accuracy of parameter estimation, but it is unlikely that they will significantly change our conclusions about the ideal choice of methodology.
@@
An important caveat to our exploration of the idealised approach to fitting is the consideration that the models themselves are also inherently uncertain. While the choice of fitting strategy, e.g. subtraction, may bias our results, the bias is (at least in most of our simulated scenarios) smaller than the systematic uncertainty in the kilonova models themselves due to uncertain underlying physics (Sarin & Rosswog, 2024; Brethauer et al., 2024). Therefore, we expect broad (order of magnitude estimates) of kilonova properties from subtraction-based fitting strategies (e.g., Rastinejad et al., 2022; Rastinejad et al., 2024) are likely robust. However, the bias is likely more of a concern for the relatively more certain afterglow parameters. We leave a detailed exploration with real observations for future work. We note that regardless of the size of bias a fit based on individual component fitting or subtraction creates, it is not the ideal quantity one desires when estimating parameters through Bayesian inference; a marginalised posterior distribution, dependent only on the observed data. Instead, these alternative approaches implicitly condition the posterior on the choice and parameters of a best-fitting afterglow model or some other underlying assumption. This is not ideal when what we truly desire is a posterior marginalising over every other degree of freedom.
@@ Acknowledgements
W. Wallace would like to thank Nordita for hosting her visit for the duration of this project, and acknowledges support for the mobility from a Turing Scheme grant awarded through the University of Bath, and support from the National Aeronautics and Space Administration (NASA) under award number 80GSFC24M0006. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of NASA. N. Sarin acknowledges support from the Knut and Alice Wallenberg Foundation through the "Gravity Meets Light" project and by the research environment grant “Gravitational Radiation and Electromagnetic Astrophysical Transients” (GREAT) funded by the Swedish Research Council (VR) under Dnr 2016-06012 and the support through the Nordita Fellowship. Nordita is supported in part by NordForsk.